Brady Construction Company contracted to build an apartment complex for a price of $6,700,000. Construction began in 2021 and was completed in 2023. The following is a series of independent situations, numbered 1 through 6, involving differing costs for the project. All costs are stated in thousands of dollars.

Estimated Costs to Complete
Costs Incurred During Year (As of the End of the Year)
Situation 2021 2022 2023 2021 2022 2023
1 1,670 2,640 1,410 4,050 1,410
2 1,670 1,410 3,080 4,050 3,080
3 1,670 2,640 2,960 4,050 2,860
4 670 3,170 1,340 4,690 960
5 670 3,170 2,490 4,690 2,860
6 670 3,170 3,500 6,255 3,230


Required:
Complete the following table. (Do not round intermediate calculations. Enter answers in dollars. Round your final answers to the nearest whole dollar. Negative amounts should be indicated by a minus sign.)
Answer #1

Working notes

Situation

Cost incurred during
the year (A)

Estimated costs
to complete (B)

Total costs
( C) = (A +B)

% of completion
D = A/C

Contract revenue
E

Gross margin
= (E ) - ( C)

Gross profit
recognized
(G ) = F X D

1

                          1,670,000

            4,050,000

         5,720,000

29.20%

               6,700,000

                  980,000

         286,119

2

                          1,670,000

            4,050,000

         5,720,000

29.20%

               6,700,000

                  980,000

         286,119

3

                          1,670,000

            4,050,000

         5,720,000

29.20%

               6,700,000

                  980,000

         286,119

4

                              670,000

            4,690,000

         5,360,000

12.50%

               6,700,000

              1,340,000

         167,500

5

                              670,000

            4,690,000

         5,360,000

12.50%

               6,700,000

              1,340,000

         167,500

6

                              670,000

            6,255,000

         6,925,000

9.68%

               6,700,000

                (225,000)

         (21,769)

For year 2022

Situation

Cost incurred during
the year (A)

Estimated costs
to complete (B)

Total costs
( C) = (A +B)

% of completion
D = A/C

Contract revenue
E

Gross margin
= (E ) - ( C)

Gross profit
recognized
(G ) = F X D

Gross profit
already
recognized (H)

Gross profit
recognized
I = G- H

1

                          4,310,000

            1,410,000

         5,720,000

75.35%

               6,700,000

                  980,000

         738,427

               286,119

             452,308

2

                          3,080,000

            3,080,000

         6,160,000

50.00%

               6,700,000

                  540,000

         270,000

               286,119

             (16,119)

3

                          4,310,000

            2,860,000

         7,170,000

60.11%

               6,700,000

                (470,000)

      (282,524)

               286,119

           (568,643)

4

                          3,840,000

                960,000

         4,800,000

80.00%

               6,700,000

              1,900,000

     1,520,000

               167,500

         1,352,500

5

                          3,840,000

            2,860,000

         6,700,000

57.31%

               6,700,000

                              -  

                    -  

               167,500

           (167,500)

6

                          3,840,000

            3,230,000

         7,070,000

54.31%

               6,700,000

                (370,000)

      (200,962)

               (21,769)

           (179,193)

*1670+2640

For last year

Situation

Cost incurred during
the year (A)

Contract revenue
( B)

Gross margin to be recognized
( C) = B - A

Gross profit
recognized in 2021
D

Gross profit
recognized in 2022
E

Total Gross
profit
recognized till date F = D + E

Gross profit
recognized in 2023
G = C - F

1

                          5,720,000

            6,700,000

             980,000

                      286,119

                   452,308

                  738,427

         241,573

2

                          6,160,000

            6,700,000

             540,000

                      286,119

                   (16,119)

                  270,000

         270,000

3

                          7,270,000

            6,700,000

          (570,000)

                      286,119

                 (568,643)

                (282,524)

      (287,476)

4

                          5,180,000

            6,700,000

         1,520,000

                      167,500

               1,352,500

              1,520,000

                    -  

5

                          6,330,000

            6,700,000

             370,000

                      167,500

                 (167,500)

                              -  

         370,000

6

                          7,340,000

            6,700,000

          (640,000)

                      (21,769)

                 (179,193)

                (200,962)

      (439,038)

The requirement

Revenue recognized over time

Revenue recognized upon completion

Situation

                                  2,021

                   2,022

                 2,023

2021

                        2,022

                       2,023

1

                              286,119

                452,308

             241,573

0

                               -  

                  980,000

2

                              286,119

                (16,119)

             270,000

0

                               -  

                  540,000

3

                              286,119

             (568,643)

          (287,476)

0

                 (179,193)

                (390,807)

4

                              167,500

            1,352,500

                        -  

0

              1,520,000

5

                              167,500

             (167,500)

             370,000

0

                  370,000

6

                              (21,769)

             (179,193)

          (439,038)

                      (21,769)

                 (179,193)

                (439,038)